What is the average speed travelled if I cover 8.2 kms in 18 minutes?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

What is the average speed travelled if I cover 8.2 kms in 18 minutes?
---
rate = distance/time
---
Ans: rate = 8.2 km/18 min = 0.4556 km/min
===================
Cheers,
Stan H.
==============

Answer by mananth(16949)

speed = distance /time
distance =8.2 km= 8 2/10 = 8 1/5 = 41/5
time = 18 minutes = 18/60= 3/10 hours
speed = (41/5)/(3/10)
speed = 41/5 * 10/3
speed = 82/3
=27.33 km/h